# 1]FAIR at Meta

### v1

@@ Abstract
Self-attention in transformer models is an incremental associative memory that maps key vectors to value vectors. One way to speed up self-attention is to employ GPU-compliant vector search algorithms, yet the standard partitioning methods yield poor results in this context, because (1) keys and queries follow different distributions and (2) the effect of RoPE positional encoding.
In this paper, we introduce Saap (Self-Attention with Asymmetric Partitions), which overcomes these problems. It is an asymmetrical indexing technique that employs distinct partitions for keys and queries, thereby approximating self-attention with a data-adaptive sparsity pattern.
It works on pretrained language models without finetuning, as it only requires to train (offline) a small query classifier. On a long context Llama 3.1-8b model, with sequences ranging from 100k to 500k tokens, our method typically reduces by a factor 20 the fraction of memory that needs to be looked-up, which translates to a time saving of 60% when compared to FlashAttention-v2.
@@ 1 Introduction
Current transformer-based large language models (LLMs)  (Devlin et al., 2019; Chowdhery et al., 2022; Brown et al., 2020; Vaswani et al., 2017) implement contextual memory under the form of self-attention: The model attends to previously processed tokens in order to predict the next one. To avoid recomputing the previous keys and values, a key-value (KV) cache is stored and reused, allowing for faster next-token prediction.
The KV cache is memory intensive: on a Llama 3-8B model Dubey et al. (2024), a single token involves a cache of size 128 kiB – about 10000 times larger than storing the token itself. Therefore storing the KV cache becomes increasingly memory-demanding as the context length increases. This is partially addressed by quantization to compress key and value vectors (Shi et al., 2024). However compression involves a trade-off, as the generation quality can suffer from information loss. Long-context self-attention is also computationally expensive, as it requires matching queries one by one to the entire KV cache, which becomes increasingly problematic as the context grows.
@@
Instead, we advocate approximating nearest-neighbor search by partitioning the keys into buckets. Practically, this enables a sparse, data-dependent, self-attention: At search time, we leverage only a small subset of buckets for each query, reducing memory access, and therefore time complexity. Multiple choices exist for the partition: random projections, as in original Locality Sensitive Hashing methods Datar et al. (2004), or data-dependent partitions, like k-means. The latter offer better trade-offs in practice on typical approximate neighbor search tasks  Paulevé et al. (2010). When using this approach for retrieving keys, the slow k-means training is performed offline on a held-out training set of keys. After this offline phase, and during the pre-fill phase with a new context and therefore new keys, performing the assignment to a given bucket of the partition is fast, as it amounts to finding the nearest centroid for each new key.
However, leveraging approximate k-nearest neighbor (ANN) methods to improve the KV cache associative faces several limitations and challenges:
@@
    top-$k$ search is harmfully restrictive for queries for which there is significant useful information in the keys that are not retrieved.
@@
    current brute-force attention implementations, such as FlashAttention-v2 (Dao, 2024) are highly tuned for the training and inference hardware. In contrast, many ANN algorithms are not GPU-compliant.
To address these limitations, we propose a novel approach called Sparse Attention with Asymmetric Partitions (Saap), which casts the partitioning of key and query vectors as a classification task: the bucket membership is predicted separately for keys and query vectors. In addition to this approach, we make the following contributions:
@@
    We quantitatively evaluate and analyze the out-of-distribution nature of query-key matching in Section 3;
@@
    We mitigate the loss of information that occurs when one limits the attention to top-$k$ keys. Instead we leverage the partition structure (Section 4.3), which leads to a variable size sparsity pattern, and show that this choice improves the self-attention approximation ;
@@
    We propose a way to efficiently run the partition-based search on GPUs, based on an optimized batched implementation (Section 4.4).
Our Saap approach reduces the resources needed to compute the self-attention of a model without finetuning it, in a hardware-compatible manner. Saap decreases the time taken by self-attention kernels by more than 60% compared to FlashAttention-v2 without compromising the generation quality, as shown by experiments in Section 5.
@@ 2 Related work
Table 1: Saap v.s. comparable KV-indexing methods: fast implementations like flashattention-v2 corresponds to the baseline. The others are indexing methods that induce sparse-attention.
@@
In this section, we review related work on accelerating attention for autoregressive transformers. In Table 1, we compare our method with a selection of the most relevant approaches along with their characteristics.
KV cache compression.
@@
Multiple techniques address the large size of the KV cache for long sequences. The approaches for this are often based on vector compression and pruning, similar to neural network pruning (LeCun et al., 1989) and compression (Han et al., 2015). The most straightforward way is to adopt low-precision computation, or more generally, variants of scalar quantization, to the KV cache. This is effective until the number of bits per weight reaches a minimum (around 4 bits), below which the compression severely degrades the accuracy (Li et al., 2024a), even when the compression is accounted for during training (Adepu et al., 2024) using quantization-aware training (Hubara et al., 2018). The keys and values can also be compressed by reducing the dimensionality of the vectors, e.g. with Principal Component Analysis (Kang et al., 2024).
@@
Another approach to compress the KV cache is to prune the least-used keys and vectors by measuring their utilization on-the-fly (Ge et al., 2023), maintaining a cache of the most used vectors (Liu et al., 2024b), ignoring the oldest keys (Xiao et al., 2024c), or using a heuristic such as H2O (Zhang et al., 2023). The GEAR method (Kang et al., 2024) combines pruning, dimensionality reduction and vector compression. For more context on KV cache compression and compute optimization, we refer the reader to the recent overview by Shi et al. (2024).
Our Saap approach is complementary to compression and static pruning. Indeed, partition-based vector search techniques are often combined with compression, see for instance the canonical IVFADC method (Jegou et al., 2010).
Vector search data structures.
@@
Several families of vector indexing structures have been applied to KV cache indexing. In the vector search literature, the “keys” are called “database” or “reference” vectors. Tree-based indexes are inspired by one- and low-dimensional search such as the seminal KD-tree Bentley (1975), but are not effective in high dimensions. Locality Sensitive Hashing (LSH) is based on hash tables where buckets store the database vectors (Datar et al., 2004). The advantage of LSH is its fast indexing time, as it amounts to computing a limited number of dot products for each key. This advantage has led to its use in several recent works on KV cache indexing (Zandieh et al., 2023; Chen et al., 2024b).
Remark: Appendix 7 discusses LSH-based methods and shows that it performs poorly in our context: vanilla LSH is not data-adaptive, so the number of independent hash tables has to be increased significantly to achieve a decent retrieval accuracy. This consumes a lot of memory and overall does not provide a significant gain with pretrained models.
@@
Graph-based indices store database vectors as nodes and searching hops from node to node (Dong et al., 2011; Malkov & Yashunin, 2018; Fu et al., 2017; Chen et al., 2024a; Jayaram Subramanya et al., 2019; Ootomo et al., 2024). Graph-based search is efficient because many routing decisions are built into the graph during its construction. Therefore, RetrievalAttention (Liu et al., 2024a) applies graph-based index RoarGraph Chen et al. (2024a) to the KV cache. The downsides are that the index building is slow and the graph is bulky to store, as each node is linked to dozens of neighbors. The consequence is that graph indexing of a KV cache is attractive only if the same KV cache is reused, e.g. because multiple prompts use the same context.
Vector search based on partitions.
The reference dataset is divided into $C$ buckets. Partitions are defined by a function that assigns each vector to one bucket. The classical variant employs an inverted file (IVF) Witten et al. (1999). When the partitioning is based on k-means, as advocated by Paulevé et al. (2010), the usage is as follows:
  • At training time, k-means clustering to the $N$ database vectors provides a set of $C$ centroids.
  • At indexing time, the assignment of a new vector amounts to finding its nearest centroid. Ideally, the IVF structure stores the vectors associated to one bucket in contiguous memory.
  • At search time, the query vector is likewise assigned to the nearest centroid and all the database vectors found in the bucket are compared with the query vector to find the top-$k$ nearest neighbors. A straightforward extension is to visit several nearest buckets instead of the single nearest one (Lv et al., 2007).
The NeuralLSH method Dong et al. (2019) predicts the bucket corresponding to the queries and database vectors alike using a neural network. In contrast with our approach it does not attempt to handle out-of-distribution data.
@@
Vector search tooling such as Faiss has been used for KV cache lookups in the Unlimiformer method (Bertsch et al., 2023), in the setting of an encoder-decoder LLM where the entire decoder shares a single cache, independently of the layer. This factorization is not possible in decoder-only LLM architectures, where each layer and each head needs to store the (key, value) pairs independently.
Out-of-distribution (OOD) search.
@@
By default, vector search assumes query and database vectors follow the same distribution. OOD search is needed when it is not the case, for example if the query and database vectors come from different modalities as with text-to-image search (Simhadri et al., 2024) or user-to-item search in recommendation systems (Paterek, 2007). The effects of OOD search on partition-based indexes, as identified by Jaiswal et al. (2022) and Chen et al. (2024a), are that (1) the nearest database vectors to the query are spread over many buckets, and (2) the query assignment does not visit the correct buckets. The authors adapt graph-based search to OOD search by taking into account how the graph navigation must be re-routed, based on a training set of query vectors.
@@
KV cache indexing has challenging OOD characteristics that we discuss in Section 3, and that our method Saap addresses in a principled way in Section 4. Note, Dedrift (Baranchuk et al., 2023) addresses temporal drift in vector databases for partition-based indexes by re-training mid-way during key ingestion. OUr Saap approach mitigates the temporal biases in a simpler way, see Section 3.4.
@@ 3 Preliminary discussion
Figure 1: Illustration of the OOD statistics between keys and queries in the attention computation. Left: first two PCA dimensions of the keys and queries (the big dots are the attention sinks for keys and queries). Middle: distribution of dot products between a random subset of keys and queries. Right: after clustering of keys, probability of assignment to each cluster, for keys (x-axis) and queries (y-axis).
@@
3.1 Attention as a vector search application
Considering the self-attention computation, the main differences with respect to classical approximate nearest-neighbor (ANN) settings are (i) there is not a hard cut-off for relevant results, but a soft weighting of results (softmax); (ii) a maximum inner product search problem instead of the more common Euclidean or cosine-similarity search; and (iii) the keys and queries are produced by different computations and thus reside in different embedding spaces. We approximate exhaustive attention using a hard cut-off on the maximum number of retrieved keys, to speed up the computation.
3.2 Out-of-distribution setting
Classical vector search methods are designed to find vectors that are sampled i.i.d. from the same distribution. We examine two types of OOD behavior: between keys and queries, and between training and testing. Prior work has quantified the difference between key and query distributions by comparing histograms of query-to-key distances v.s. key-to-key distances (Jaiswal et al., 2022; Chen et al., 2024a).
Saap is based on a partition of the keys. For a query vector, the $\ell$ most promising buckets are visited and the dot product between the query and all keys from the visited buckets is computed. One important aspect that plays into OOD behavior for Saap, is the degree to which the buckets of the $N$ key vectors are balanced in terms of membership. If the buckets are all of the same size and equally likely to be visited, then the number of dot products to perform is $\ell N/C$. The search time is proportional to this.
3.3 Key-query OOD
Keys and queries are computed using distinct linear layers on top of shared vectors. The distribution mismatch between keys and queries is observed with basic statistics of a random sample of vectors (here head 28 of layer 17 of a Llama 3 model). Figure 1(left) shows the first principal components of keys and queries: their embedding spaces are clearly distinct. As the two point clouds are on opposite sides of the origin, the $\langle k,q\rangle$ dot products are almost always negative (Figure 1, middle). Note, the dot products with the first key (i.e. the “attention sink” (Xiao et al., 2024c)) and the most recent keys, are statistically higher than the other keys.
From now on, we apply a spherical k-means clustering with $C=1024$ centroids to the keys (Douze et al., 2024). Figure 1 (right) shows that the probability of assignment to a given centroid are very different between keys and queries. More specifically, some clusters “capture” a significant fraction of the queries (up to 10%). Motivated by this, we re-design an alternative assignment operator in Section 4.
@@ 3.3 Key-query OOD
3.4 Prompt and temporal ODD
OOD problems occur for all ML models that assume i.i.d. training and test data. Here we examine the divergence of the distribution at training time compared to at test time when a prompt is used, focusing entirely on the key vector distribution. We measure the variance of the assignment of 100k vectors sampled in different ways, and repeat the experiment 10 times: higher variance indicates stronger OOD behavior. Figure 2 shows the minimal variance occurs when vectors are sampled uniformly on the training set. When we measure it on 100k random uniform vectors from data that is disjoint from the training set, the variance increases. We sampled 100k vectors from each of 10 different prompts and found that the variance changes significantly depending on the prompt: this is prompt OOD. We then sample 100k vectors across prompts, but from 10 narrow timestamp ranges: here the deviation with respect to the training distribution is the highest, showing strong temporal bias.
Removing temporal bias.
The rope temporal transformation (Su et al., 2023), when applied to all keys and queries, is likely a major contributor to their temporal bias. Therefore, we repeated our analysis on keys without the rope transformation and found that, indeed the time bias is reduced (blue dots in Figure 2). We then applied the same de-rope operation to queries where the rope component has been removed, yielding the last plot in Figure 2. Given the positive effect of de-roping, we use it for all models trained for Saap.
3.5 Per-head attention span
Figure 1 (middle) shows that the first key of any sequence, the so-called attention sink (Xiao et al., 2024c), is an outlier in the data distribution and always has a large dot product with the query vectors. Similarly, the most recent keys have higher dot products with the queries. Therefore, it is typical in attention approximations to retrieve these keys exactly. Approximate and exact retrieval results are then combined to form a final attention result. In our case, we use the “1+2047” setting, which means that the first token and the most recent 2047 tokens are attended exhaustively. This context is long enough to cover many short-context tasks.
Figure 3 shows the fraction of attention that comes from the top 256 keys (out of $10^{5}$) for each query (excluding the attention sink and recent tokens). For most heads, the top-256 cover most of the attention weight. Layer 0 is an exception, all its heads are more uniformly spread over keys. Therefore, in the following, we perform full attention for the layer 0, similarly to Tang et al. (2024); Xiao et al. (2024c).

### v2

# Inference-time sparse attention with asymmetric indexing

@@ Abstract
Self-attention in transformer models is an incremental associative memory that maps key vectors to value vectors. One way to speed up self-attention is to employ GPU-compatible vector search algorithms based on standard partitioning methods such as k-means. However, such partitioning methods yield poor results in this context because (1) the keys and queries follow different distributions, and (2) the RoPE positional encoding hinders the bucket assignment.
This paper introduces Saap (Self-Attention with Asymmetric Partitions), which overcomes these problems. It is an asymmetrical indexing technique that employs distinct partitions for keys and queries, thereby approximating self-attention with a data-adaptive sparsity pattern. It works on pretrained language models and only requires to train (offline) a small query classifier. On a long context Llama 3.1-8b model, with sequences ranging from 100k to 500k tokens, Saap typically reduces by a factor of 20 the fraction of memory that needs to be looked-up, which translates to a time saving of 60% when compared to FlashAttention-v2.
@@ 1 Introduction
Current transformer-based large language models (LLMs)  (Vaswani et al., 2017; Brown et al., 2020) implement contextual memory in the form of self-attention: The model attends to previously processed tokens in order to predict the next one. To avoid recomputing the previous keys and values, a key-value (KV) cache is stored and reused, allowing for faster next-token prediction. This KV cache is memory intensive: on a Llama 3-8B model (Dubey et al., 2024), a single token involves a cache of size 128 kiB – about 10000 times larger than storing the token itself. Therefore, this cache becomes increasingly memory-demanding as the context length increases. This problem is partially addressed by quantization to compress key and value vectors (Shi et al., 2024) but comes at a cost, as the generation quality eventually degrades below 4 bits per weight. Additionally, long-context self-attention requires matching queries one by one to the entire KV cache, which becomes computationally prohibitive as the context grows.
@@
Instead, we advocate approximating nearest-neighbor search by partitioning the keys into buckets. Practically, this enables sparse, data-dependent, self-attention: at search time, we leverage only a small subset of buckets for each query, reducing memory access, and therefore time complexity. Multiple choices exist for the partition: random projections, as in the original locality-sensitive hashing methods (Datar et al., 2004), or data-dependent partitions, such as k-means. The latter offers better trade-offs in practice in typical approximate neighbor search tasks  (Paulevé et al., 2010). In this approach for retrieving keys, the slow k-means training is performed offline on a held-out training set of keys. After this offline phase, and during the pre-fill phase with a new context (and therefore new keys), performing the assignment to a given bucket of the partition is fast, since it amounts to finding the nearest centroid for each new key.
However, in this context, approximate k-nearest neighbor (ANN) methods to improve KV cache efficiency face several limitations and challenges:
@@
    retrieving the top-$k$ keys only is restrictive for queries where useful information is associated to keys beyond the top-$k$.
@@
    current brute-force attention implementations, such as FlashAttention-v2 (Dao, 2024) are highly tuned for the training and inference hardware. In contrast, many ANN algorithms are not GPU-compatible.
To address these limitations, we propose a novel approach called Sparse Attention with Asymmetric Partitions (Saap), that casts the partitioning of key and query vectors as a classification task: the bucket membership is predicted separately for keys and query vectors. In addition to this approach, we make the following contributions:
@@
    We quantitatively analyze the out-of-distribution nature of query-key matching (Section 3) and introduce an asymmetrical assignment strategy for keys and queries (Section 4);
@@
    We mitigate the loss of information that occurs when the attention is limited to top-$k$ keys. Instead we leverage the partition structure (Section 4), which leads to a variable size sparsity pattern, and show that this choice improves the self-attention approximation;
@@
    We propose a method for efficiently running the partition-based search on GPUs, based on an optimized batched implementation (Section 5).
Saap reduces the resources needed to compute the self-attention of a model without finetuning it, in a hardware-compatible manner. Compared to the state-of-the-art FlashAttention-v2 baseline, it decreases the time taken by self-attention kernels by more than 60%, without compromising the generation quality, see Section 6.
@@ 2 Related work
Table 1: Saap v.s. comparable KV-indexing methods: fast implementations like flashattention-v2 correspond to the full attention baseline. The other columns correspond to indexing methods that induce sparse attention.
@@
This section reviews related work on accelerating attention for autoregressive transformers. Table 1 compares our method with a selection of the most relevant approaches along with their characteristics.
@@
KV cache compression.   Multiple techniques address the large size of the KV cache for long sequences. Some approaches rely on vector compression and pruning, similar to neural network pruning (LeCun et al., 1989) and compression (Han et al., 2015). The most straightforward way is to adopt low-precision computation variants of scalar quantization for the KV cache. This is effective until the number of bits per weight becomes detrimental (around 4 bits) and the compression severely degrades the accuracy (Li et al., 2024a), even when the compression is accounted for during training (Adepu et al., 2024) with quantization-aware training (Hubara et al., 2018). The keys and values can also be compressed by dimensionality reduction, e.g. with Principal Component Analysis (Kang et al., 2024).
@@
Another approach to compress the KV cache is to prune the least-used keys and vectors by measuring their utilization on-the-fly (Ge et al., 2023), maintaining a cache of the most used vectors (Liu et al., 2024b), ignoring the oldest keys (Xiao et al., 2024b), or using a heuristic such as H2O (Zhang et al., 2023). The GEAR method (Kang et al., 2024) combines pruning, dimensionality reduction and vector compression. For more context on KV cache compression and compute optimization, we refer the reader to the recent overview by Shi et al. (2024). Our Saap approach is complementary to compression and static pruning. Indeed, partition-based vector search techniques are often combined with compression, for instance, the canonical IVFADC method (Jegou et al., 2010).
@@
Vector search data structures.   Several families of vector indexing structures have been applied to KV cache indexing. In the vector search literature, the “keys” are called “database” or “reference” vectors. Tree-based indexes are inspired by one- and low-dimensional search such as the seminal KD-tree (Bentley, 1975), but are not effective in high dimensions. Locality Sensitive Hashing (LSH) is based on hash tables where buckets store the database vectors (Datar et al., 2004). The advantage of LSH is its fast indexing time, as it amounts to computing a limited number of dot products for each key. This advantage has led to its use in several recent works on KV cache indexing (Zandieh et al., 2023; Chen et al., 2024b). In Appendix B, we discuss LSH-based methods and show that it performs poorly for KV cache acceleration: vanilla LSH is not data-adaptive, so the number of independent hash tables has to be increased significantly to achieve a decent retrieval accuracy. This consumes a lot of memory and overall does not provide a significant gain with pretrained models.
@@
Graph-based indices store database vectors as nodes and searching hops from node to node (Dong et al., 2011; Malkov and Yashunin, 2018; Fu et al., 2017; Jayaram Subramanya et al., 2019). Graph-based search is efficient because routing decisions are built into the graph during its construction. Therefore, RetrievalAttention (Liu et al., 2024a) applies graph-based index RoarGraph (Chen et al., 2024a) to the KV cache. The downsides are that the index building is slow and the graph is bulky to store, as each node is linked to dozens of neighbors. As a result, graph indexing of a KV cache is attractive only if the same KV cache is reused, e.g. because multiple prompts use the same context.
Vector search based on partitions.   With this approach the reference dataset is partitioned into $C$ buckets. Partitions are defined by a function that assigns each vector to one bucket. The classical variant employs an inverted file (IVF) (Witten et al., 1999). When the partitioning is based on k-means, as advocated by Paulevé et al. (2010). The usage is as follows: (1) At training time, k-means clustering of the $N$ database vectors provides a set of $C$ centroids. (2) At indexing time, the assignment of a new vector amounts to finding its nearest centroid. Ideally, the IVF structure stores the vectors associated to one bucket in contiguous memory. (3) At search time, the query vector is likewise assigned to the nearest centroid and all the database vectors found in the bucket are compared with the query vector to find the top-$k$ nearest neighbors. A straightforward extension is to visit several nearest buckets instead of the single nearest one (Lv et al., 2007). The NeuralLSH method (Dong et al., 2019) predicts the bucket corresponding to the queries and database vectors alike using a neural network. Unlike our approach, it does not handle out-of-distribution data.
@@
Vector search tooling such as Faiss has been used for KV cache lookups in the Unlimiformer method (Bertsch et al., 2023), in the setting of an encoder-decoder LLM where the entire decoder shares a single cache, independently of the layer. This factorization is not possible in decoder-only LLM architectures, where each layer and each head need to store the (key, value) pairs independently.
@@
Out-of-distribution (OOD) search.   By default, vector search assumes query and database vectors follow the same distribution. OOD search is needed when it is not the case, for example, if the query and database vectors come from different modalities as with text-to-image search (Simhadri et al., 2024) or user-to-item search in recommendation systems (Paterek, 2007). The effects of OOD search on partition-based indexes, as identified by Jaiswal et al. (2022) and Chen et al. (2024a), are: (1) the nearest database vectors to the query are spread over many buckets, and (2) the query assignment does not visit the correct buckets. The authors adapt graph-based search to OOD search by taking into account how the graph navigation must be re-routed, based on a training set of query vectors.
@@
KV cache indexing has challenging OOD characteristics that we discuss in Section 3, and that our method Saap addresses in a principled way in Section 4. Dedrift (Baranchuk et al., 2023) addresses temporal drift in vector databases for partition-based indexes by re-training mid-way during key ingestion. Our Saap approach mitigates the temporal biases in a simpler way, see Section 3.
Figure 1: Illustration of the OOD statistics between keys and queries in the self-attention. Left: first two PCA dimensions of keys and queries (the big dots are the attention sinks for keys and queries). Middle: distribution of dot products between random subsets of keys and queries. Right: after clustering keys, probability of assignment to each cluster for keys (x-axis) and queries (y-axis).
@@ 3 Preliminary discussion
Attention as a vector search application   Considering the self-attention computation, the main differences with respect to classical approximate nearest-neighbor (ANN) settings are: (i) there is not a hard cut-off for relevant results, but a soft weighting of results using softmax; (ii) it is a maximum inner product search problem instead of the more common Euclidean or cosine-similarity search; and (iii) the keys and queries are produced by different computations, thus, they reside in different embedding spaces. To accelerate the computation, we approximate exhaustive attention using a hard cut-off on the maximum number of retrieved keys.
Out-of-distribution setting and partitioning of keys   Classical vector search methods are designed to find vectors that are sampled i.i.d. from the same distribution. We examine two types of OOD behavior: (i) between keys and queries, and (ii) between training and testing. Prior work has quantified the difference between key and query distributions by comparing histograms of query-to-key distances v.s. key-to-key distances (Jaiswal et al., 2022; Chen et al., 2024a). Saap is based on a partition of the keys. For a query vector, the $\ell$ most promising buckets are visited and the dot product between the query and all keys from the visited buckets is computed. One important aspect that plays into OOD behavior for Saap, is the degree to which the buckets of the $N$ key vectors are balanced in terms of membership. If the buckets are all of the same size and equally likely to be visited, the number of dot products to perform is $\ell N/C$. The search time is proportional to this.
Key-query OOD   Keys and queries are computed using distinct linear layers on top of shared vectors. The distribution mismatch between keys and queries is observed with basic statistics of a random sample of vectors (here head 28 of layer 17 of a Llama 3 model). Figure 1 (left) shows the first principal components of keys and queries: their embedding spaces are clearly distinct. As the two point clouds are on opposite sides of the origin, the $\langle k,q\rangle$ dot products are almost always negative (Figure 1, middle). The dot products with the first key (i.e. the “attention sink” (Xiao et al., 2024b)) and the most recent keys, are statistically higher than the other keys.
[…]
@@
Prompt and temporal OOD   OOD problems occur for all ML models that assume i.i.d. training and test data. Here we examine the divergence of the distribution at training time compared to at test time when a prompt is used, focusing entirely on the key vector distribution. We measure the variance of the assignment of 100k vectors sampled in different ways, and repeat the experiment 10 times: higher variance indicates stronger OOD behavior. Figure 2 shows that the minimal variance occurs when vectors are sampled uniformly from the training set. When we measure it on 100k random uniform vectors from data that is disjoint from the training set, the variance increases. We sampled 100k vectors from each of 10 different prompts and found that the variance changes significantly depending on the prompt: this is prompt OOD. We then sample 100k vectors across prompts, but from 10 narrow timestamp ranges: here the deviation with respect to the training distribution is the highest, showing strong temporal bias.
Removing temporal bias.   The rope temporal transformation (Su et al., 2023) is likely a major contributor to their temporal bias when applied to all keys and queries. Therefore, we repeated our analysis on keys without the rope transformation and found that indeed the time bias is reduced (blue dots in Figure 2). We then applied the same de-rope operation to queries where the rope component has been removed, yielding the last plot in Figure 2. Given the positive effect of de-roping, we use it for all models trained for Saap.
Per-head attention span   Figure 1 (middle) shows that the first key of any sequence, the so-called attention sink (Xiao et al., 2024b), is an outlier in the data distribution and always has a large dot product with the query vectors. Similarly, the most recent keys have higher dot products with the queries. Therefore, it is typical in attention approximations to retrieve these keys exactly. Approximate and exact retrieval results are then combined to form a final attention result. In our case, we use the “1+2047” setting, which means that the first token and the most recent 2047 tokens are attended exhaustively. This context is long enough to have good performance on many short-context tasks.
Figure 3 shows the fraction of attention that comes from the top 256 keys (out of $10^{5}$) for each query (excluding the attention sink and recent tokens). For most heads, the top-256 cover most of the attention weight. Layer 0 is an exception, all its heads are more uniformly spread over keys. Therefore, in the following, we perform full attention for the layer 0, similarly to the work by Tang et al. (2024); Xiao et al. (2024b).